OpenStudy (anonymous):

need help with evaluating sum notations:

OpenStudy (anonymous):

if \[\sum_{i=1}^{10}a_{i}=14 \] and \[ \sum_{i=1}^{20}a_{i}=5\] Evaluate the following sums: \[a) \sum_{i=1}^{10}13-a_{i}\] There's a few more, but I think i'll be able to understand the more complex ones if I get this one.

OpenStudy (anonymous):

\[ \sum_{i=1}^{10}13-a_{i} = \sum_{i=1}^{10}13- \sum_{i=1}^{10}a_{i} \]

OpenStudy (anonymous):

\[ \sum_{i=1}^{10}13-a_{i} = \sum_{i=1}^{10}13- \sum_{i=1}^{10}a_{i} = 13(10)-14 \]

OpenStudy (anonymous):

@dw3 Make sense? If so, medal?

OpenStudy (anonymous):

Got it. Cheers.